Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-10-10

Observers: S.Moran, T. Reynolds
Support Team: J. Harmanen, H. Kuncarayakti
TAT: Régis Cartier, Michel Dennefeld

Calibrations: 
Flats and arcs for gr11,13,16,18,20 + biases

Weather: 
Start - Thick clouds (closed)
Middle - Thick clouds (closed)
End - Thick clouds (closed)

UT       Target       Type      Setup          Exp         airmass  DIMM  note
-----------------------------------------------------------------------------------------
                                             
EFOSC         
Could not open due to thick clouds all night.
Comments